OpenStudy (anonymous):

How do I write a slope intercept form with these points. If a line passes through (-3,-1), parallel to the line that passes through (3, 3) and (0, 6)? Please help!!!!

OpenStudy (anonymous):

slope of (3, 3) and (0, 6) is the same slope with the line passing through (-3,-1) because they are parallel 1. find the slope of (3, 3) and (0, 6), using the slope formula 2.write the slope intercept from(y=mx+b) using the equation of the line[(y-y1)=m(x-x1)]

OpenStudy (anonymous):

Thnx so much!! Is it with the same thing to the other question I asked?

OpenStudy (anonymous):

what other question?

OpenStudy (anonymous):

Hold on to that question...I got -x what should I do?

OpenStudy (anonymous):

i'm not sure what you mean..can you show me your solution?

OpenStudy (anonymous):

I got y= -1x +6

OpenStudy (anonymous):

its ok if it looks like that

OpenStudy (anonymous):

yeah, it's ok.. y= -x+6